Let 𝑓(π‘₯) = 1 3 π‘₯ 3 + π‘₯ 2 βˆ’ 15π‘₯ βˆ’ 9 . Use detailed sign tables in answering the following questions. (a) Find the intervals in which 𝑓 is increasing or decreasing. (b) Find the intervals in which the graph of 𝑦 = 𝑓(π‘₯) is concave upwards or downwards.Β 



1
Expert's answer
2021-04-29T14:43:14-0400

"f(x)=13x^3+x^2-15x-9"


a)

"f'(x)=39x^2+2x-15"


"f'(x)=0" when "x=\frac{-2\pm\sqrt{4+4*15*39}}{2*39}=\frac{-1\pm\sqrt{586}}{39}"


"f(x)" is increasing when "f'(x)>0" or when

"-\infty<x<\frac{-1-\sqrt{586}}{39}"

and when

"\frac{-1+\sqrt{586}}{39}<x<+\infty"


"f(x)" is decreasing when "f'(x)<0" or when

"\frac{-1-\sqrt{586}}{39}<x<\frac{-1+\sqrt{586}}{39}"


b)

"f''(x)=78x+2"


"f''(x)=0" when "x=-\frac{1}{39}"


When "x<-\frac{1}{39}" , "f''(x)<0" , "f(x)" is concave upwards.


When "x>-\frac{1}{39}" , "f''(x)>0" , "f(x)" is concave downwards.